OpenStudy (anonymous):

I am desprate for math help!! a^-2 = 1/36 <---- Really confused What is the value of the expression -x + y when x = 1.2 and y = -10.7 <--- Got halfway through the problem then forgot what to do/

OpenStudy (the_fizicx99):

What's the question, a^-2 = 1/36 or What is the value of the expression -x + y when x = 1.2 and y = -10.7

OpenStudy (anonymous):

Both questions are by themselves

OpenStudy (the_fizicx99):

What? ._. So its 2 questions?

OpenStudy (anonymous):

Yes

OpenStudy (the_fizicx99):

For the 2nd plug in -x + y when x = 1.2 and y = -10.7 -1.2 + (-10.7) + and a - = - so its -1.2 - 10.7 = ???

OpenStudy (anonymous):

-11.9

OpenStudy (the_fizicx99):

Yup

OpenStudy (anonymous):

Thanks and the first one?

OpenStudy (the_fizicx99):

What're the options for the first one, I;m not sure how to go about it..?

OpenStudy (anonymous):

The options were A. 6 B. 4 C. -9 D -18

OpenStudy (the_fizicx99):

Is that the entire question or does it say plug in for a and solve, because you can't have a negative you would put it 1/a^2

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

It is the whole question

OpenStudy (anonymous):

I will just guess

OpenStudy (the_fizicx99):

A

OpenStudy (anonymous):

Thank you sooooooooooo much!!!!!!!!!!!

OpenStudy (the_fizicx99):

Yw